Jung's average on 2 biology quizzes is 7. what should Jung's score on the next quiz be in order to have an average of 8 for the 3 quizzes?
Found 2 solutions by mananth, greenestamps:
Answer by mananth(16946)


Jung's average on 2 biology quizzes is 7. what should Jung's score on the next quiz be in order to have an average of 8 for the 3 quizzes?

Jung's average on 2 biology quizzes is 7. Total = 2*7 =14
Required average = 8 = (14 +x)/3 where x is the score in third test
24 = x+14
24-14 =x
x=10
The score should be 10 in the third

Answer by greenestamps(13206)



The response from the other tutor shows a formal mathematical solution... which possibly is what was wanted.

If a formal algebraic solution is not required, you can solve a simple averaging problem like this one as follows:

His point total on the first two quizzes is 14; to get the desired average of 8, his point total would have to be 16. So on the first two quizzes he is 2 points BELOW where he needs to be.

That means on the next quiz, in order to reach his desired average, his score needs to be 2 points ABOVE the desired average.

ANSWER: The required score on the next quiz is 8+2 = 10.